OpenStudy (anonymous):

Three party-goers are in the corner of the ballroom having an intense argument. You walk over to settle the debate. They are discussing a function g(x). You take out your notepad and jot down their statements. Correct the reasoning of any inaccurate reasoning by the party-goers in full and complete sentences. Make sure you reference any theorems that support your justifications. o Professor McCoy: She says that 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. o Ms. Guerra: She says that 2 is a zero of g(x) because g(2) = 0. o Mr. Romano: He says that 2 is

OpenStudy (anonymous):

o Mr. Romano: He says that 2 is a zero of g(x) because synthetic division with 2 results in a remainder of 0.

OpenStudy (anonymous):

When I see "intense argument", "party-goers", and "discussing a function", I know this question isn't serious.

OpenStudy (anonymous):

It's a word problem..

OpenStudy (anonymous):

Ms. Guerra: She says that 2 is a zero of g(x) because g(2) = 0. This statement looks right.

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

an x value that when plugged into an expression turns out zero is a zero. For example... If g(x) = x^2 -4 you know that you can find x by adding 4 and taking the square root and x = 2 g(2) = 2^2 - 4 = 0 2 is a zero.

OpenStudy (anonymous):

x could = -2 also but that doesn't matter for my point.

OpenStudy (anonymous):

as far as long division and synthetic division... its been a long time. Might want to get that from someone else by bumping this.

OpenStudy (anonymous):

Thanks :)